# A Set of Tutorials for the LAMMPS Simulation Package [Article v1.0]

### v1

@@ 3.6.2 Cracking the silica
The fix nvt command is employed to control the temperature of the system. As observed from the generated images, the atoms progressively adjust to the changing box dimensions. At some point, bonds begin to break, leading to the appearance of dislocations (Fig. 37).
@@ 3.6.3 Adding water
To add the water molecules to the silica, we will employ the Monte Carlo method in the grand canonical ensemble (GCMC). In short, the system is placed into contact with a virtual reservoir of a given chemical potential $\mu$, and multiple attempts to insert water molecules at random positions are made. Each attempt is either accepted or rejected based on energy considerations. For further details, please refer to classical textbooks like Ref. 1.
Using hydrid potentials
The first particularly of our system is that it combines water and silica, which necessitates the use of two force fields: Vashishta (for $\text{SiO}_{2}$), and TIP4P (for water). Here, the TIP4P/2005 model is employed for the water [13]. Open the gcmc.lmp file, which should contain the following lines:
@@
pair_style hybrid/overlay vashishta &
lj/cut/tip4p/long OW HW OW-HW HW-OW-HW 0.1546 10
@@
Combining the two force fields, Vashishta and TIP4P/2005, is achieved using the hybrid/overlay pair style. The PPPM solver [37] is specified with the kspace command, and is used to compute the long-range Coulomb interactions associated with tip4p/long. Finally, the style for the bonds and angles of the water molecules are defined; however, these specifications are not critical since TIP4P/2005 is a rigid water model.
The water molecule template called H2O.mol must be downloaded and located next to gcmc.lmp
@@ 1 HW-OW-HW
After reading the data file and defining the h2omol molecule from the H2O.txt file, the create_atoms command is used to include three water molecules in the system. Then, add the following pair_coeff (and bond_coeff and angle_coeff) commands to gcmc.lmp:
pair_coeff * * vashishta SiO.1990.vashishta Si O NULL NULL
pair_coeff * * lj/cut/tip4p/long 0 0
pair_coeff Si OW lj/cut/tip4p/long 0.0057 4.42
pair_coeff O OW lj/cut/tip4p/long 0.0043 3.12
pair_coeff OW OW lj/cut/tip4p/long 0.008 3.1589
pair_coeff HW HW lj/cut/tip4p/long 0.0 0.0
@@
The force field Vashishta applies only to Si and O of $\text{SiO}_{2}$, and not to the OW and HW of $\text{H}_{2}\text{O}$, thanks to the NULL parameters used for atoms of types OW and HW. Pair coefficients for the lj/cut/tip4p/long potential are defined between O($\text{H}_{2}\text{O}$) and between H($\text{H}_{2}\text{O}$) atoms, as well as between O($\text{SiO}_{2}$)-O($\text{H}_{2}\text{O}$) and Si($\text{SiO}_{2}$)-O($\text{H}_{2}\text{O}$). Thus, the fluid-fluid and the fluid-solid interactions will be adressed with by the lj/cut/tip4p/long potential. The bond_coeff and angle_coeff commands set the OW-HW bond length to 0.9572 Å, and the HW-OW-HW angle to $104.52^{\circ}$, respectively [13].
@@
The number of oxygen atoms from water molecules (i.e. the number of molecules) is calculated by the nO variable. The SHAKE algorithm is used to maintain the shape of the water molecules over time [46, 47].
@@
compute_modify thermo_temp dynamic yes
compute_modify ctH2O dynamic yes
fix mynvt1 H2O nvt temp 300 300 0.1
fix_modify mynvt1 temp ctH2O
fix mynvt2 SiO nvt temp 300 300 0.1
@@
Two different thermostats are used for $\text{SiO}_{2}$ and $\text{H}_{2}\text{O}$, respectively. Using separate thermostats is usually better when the system contains two separate species, such as a solid and a liquid. It is particularly important to use two thermostats here because the number of water molecules will fluctuate with time. The compute_modify command with the dynamic yes option for water is used to specify that the number of molecules will not be constant.
@@
The tfac_insert option ensures the correct estimate for the temperature of the inserted water molecules by taking into account the internal degrees of freedom. Here, 100 insertion and deletion attemps are made every 100 steps.
Note 3.22.
[…]
@@
Running this simulation using LAMMPS, one can see that the number of molecules is increasing progressively. When using the pressure argument, LAMMPS ignores the value of the chemical potential (here $\mu=-0.5\,\text{eV}$, which corresponds roughly to ambient conditions, i.e. to a relative humidity $\text{RH}\approx 50\,\%$ [56].) The large pressure value of 100 bars was chosen to ensure that some successful insertions of molecules would occur during the short duration of this simulation.
After a few GCMC steps, the number of molecules starts increasing. Once the crack is filled with water molecules, the total number of molecules reaches a plateau (Figs. 38-39). The final number of molecules depends on the imposed pressure, temperature, and the interaction between water and silica (i.e. its hydrophilicity). Note that GCMC simulations of such dense phases are usually slow to converge due to the very low probability of successfully inserting a molecule. Here, the short simulation duration was made possible by the use of a high pressure.
@@ 3.7.1 Method 1: Free sampling
The most direct way to calculate a free energy profile is to extract the partition function from a classical (i.e. unbiased) molecular dynamics simulation, and then estimate the Gibbs free energy by using
$\Delta G=-RT\ln(p/p_{0}),$ (4)
where $\Delta G$ is the free energy difference, $R$ is the gas constant, $T$ is the temperature, $p$ is the pressure, and $p_{0}$ is a reference pressure. As an illustration, let us apply this method to a simple configuration that consists of a particles in a box in the presence of a position-dependent repulsive force that makes the center of the box a less favorable area to explore.
@@
To begin this tutorial, select «Start Tutorial 7» from the «Tutorials» menu of LAMMPS–GUI and follow the instructions. The editor should display the following content corresponding to free-energy.lmp:
@@
Here, we begin by defining variables for the Lennard-Jones interaction $\sigma$ and $\epsilon$ and for the repulsive potential $U$, which are $U_{0}$, $\delta$, and $x_{0}$ [see Eqs. (5-6) below]. The cut-off value of 3.822 was chosen to create a Weeks-Chandler-Andersen (WCA) potential, which is a truncated and purely repulsive LJ potential [63]. It was calculated as $2^{1/6}\sigma$. The potential is also shifted to be equal to 0 at the cut-off using the pair_modify command. The unit system is real, in which energy is in kcal/mol, distance in Ångströms, or time in femtosecond, has been chosen for practical reasons: the WHAM algorithm used in the second part of the tutorial automatically assumes the energy to be in kcal/mol.
@@
Let us define the simulation box and randomly add atoms by addying the following lines to free-energy.lmp:
@@
[…]
We impose the force $F(x)$ to the atoms in the simulation using the fix addforce command. Add the following lines to free-energy.lmp:
@@
Next, we combine the fix nve with a fix langevin thermostat:
@@
[…]
To ensure that the equilibration time is sufficient, we will track the evolution of the number of atoms in the central - energetically unfavorable - region, referred to as mymes, using the n_center variable:
@@
dump viz all image 50000 myimage-*.ppm type type &
@@
A dump image command was also added for system visualization.
[…]
@@
[…]
@@
Once the system is equilibrated, we will record the density profile of the atoms along the $x$-axis using the ave/chunk command. Add the following line to free-energy.lmp:
@@
The step count is reset to 0 using reset_timestep to synchronize it with the output times of fix density/number. Run the simulation using LAMMPS.
@@
Once the simulation is complete, the density profile from free-sampling.dat shows that the density in the center of the box is about two orders of magnitude lower than inside the reservoir (Fig. 44 a). Next, we plot $-RT\ln(\rho/\rho_{\mathrm{bulk}})$ (i.e. Eq. (4) where the pressure ratio $p/p_{\mathrm{bulk}}$ is replaced by the density ratio $\rho/\rho_{\mathrm{bulk}}$, assuming the system behaves as an ideal gas) and compare it with the imposed potential $U$ from Eq. (5) (Fig. 44 b). The reference density, $\rho_{\text{bulk}}=0.0009\leavevmode\nobreak\ \text{\AA{}}^{-3}$, was estimated by measuring the density of the reservoir from the raw density profiles. The agreement between the MD results and the imposed energy profile is excellent, despite some noise in the central part, where fewer data points are available due to the repulsive potential.
[…]
@@
Increasing the value of $U_{0}$ reduces the average number of atoms in the central region, making it difficult to achieve a high-resolution free energy profile. For example, running the same simulation with $U_{0}=10\epsilon$, corresponding to $U_{0}\approx 10k_{\text{B}}T$, results in no atoms exploring the central part of the simulation box during the simulation. In such a case, employing an enhanced sampling method is recommended, as done in the next section.
@@ 3.7.2 Method 2: Umbrella sampling
dump viz all image 50000 myimage-*.ppm type type &
@@
So far, our code resembles that of Method 1, except for the additional particle of type 2. Particles of types 1 and 2 are identical, with the same mass and LJ parameters. However, the particle of type 2 will also be exposed to the biasing potential $V$, which forces it to explore the central part of the box (Fig. 45).
[…]
@@
The spring command imposes the additional harmonic potential $V$ with the previously defined spring constant $k$. The center of the harmonic potential, $x_{\text{des}}$, successively takes values from $-28\,\text{\AA}$ to $28\,\text{\AA}$. For each value of $x_{\text{des}}$, an equilibration step of 40 ps is performed, followed by a step of 400 ps during which the position of the particle of type 2 along the $x$-axis, $x_{\text{ave}}$, is saved in data files named umbrella-sampling.i.dat, where $i$ ranges from 1 to 15. Run the umbrella-sampling.lmp file using LAMMPS.
Note 3.23.

### v2

@@ 3.6.2 Cracking the silica
As discussed, the fix nvt command integrates the Nosé-Hoover equations of motion to sample the NVT ensemble, which allows controlling the temperature of the system. As observed from the generated images, the atoms progressively adjust to the changing box dimensions. At some point, bonds begin to break, leading to the appearance of dislocations (Fig. 37).
Note 3.40.
Although the Nosé-Hoover equations were originally formulated to sample the NVT ensemble, using the fix nvt command does not guarantee that a simulation actually samples the NVT ensemble.
@@ 3.6.3 Adding water
To add the water molecules to the silica, we will employ the Monte Carlo method in the grand canonical ensemble (GCMC). In short, the system is placed into contact with a virtual reservoir containing pure water at a given thermodynamic state, and multiple attempts to insert water molecules at random positions are made. In the grand canonical ensemble, each attempt is either accepted or rejected based on internal energy and chemical potential, $\mu$ considerations. For further details, please refer to classical textbooks like Ref. 1.
Adapting the pair style
For this next step, we need to specify the force field used to model the interactions in the system. The TIP4P/2005 model is employed for the water [13], while no interaction within silica is defined, as it will be seen farther below. This is because the atoms of the silica will remain frozen during this part of the simulation. Only the cross-interactions between water and silica need to be defined. Open the gcmc.lmp file, which should contain the following lines:
@@
pair_style lj/cut/tip4p/long OW HW OW-HW HW-OW-HW 0.1546 10
@@
The PPPM solver [38] is specified with the kspace command, and is used to compute the long-range Coulomb interactions associated with tip4p/long. Finally, the form of the bond and angle potentials of the water molecules are defined; however, as previously discussed, these specifications are not critical since TIP4P/2005 is a rigid water model.
Note 3.41.
In practice, it is possible to use both vashishta and lj/cut/tip4p/long pair styles at the same time by employing the pair_style hybrid command. However, hybridizing force fields should be done with caution, as there is no guarantee that the resulting force field will produce meaningful results.
The water molecule template called H2O.mol must be downloaded and located next to gcmc.lmp.
@@ 1 HW-OW-HW
After reading the data file and defining the h2omol molecule from the H2O.txt file, the create_atoms command is used to include three water molecules in the system. Then, add the following pair_coeff (and bond_coeff and angle_coeff) commands to gcmc.lmp in order to set the potential parameters:
pair_coeff * * 0 0
pair_coeff Si OW 0.0057 4.42
pair_coeff O OW 0.0043 3.12
pair_coeff OW OW 0.008 3.1589
pair_coeff HW HW 0.0 0.0
@@
Pair coefficients for the lj/cut/tip4p/long pair style are defined between O($\text{H}_{2}\text{O}$) and between H($\text{H}_{2}\text{O}$) atoms, as well as between O($\text{SiO}_{2}$)-O($\text{H}_{2}\text{O}$) and Si($\text{SiO}_{2}$)-O($\text{H}_{2}\text{O}$). Thus, the fluid-fluid and the fluid-solid interactions will be adressed with by the lj/cut/tip4p/long potential. The bond_coeff and angle_coeff commands set the OW-HW bond length to 0.9572 Å, and the HW-OW-HW angle to $104.52^{\circ}$, respectively [13].
Note 3.42.
The pair coefficients for interactions between Si($\text{SiO}_{2}$) and O($\text{SiO}_{2}$) are set by the first command, pair_coeff * * 0 0, which effectively means that they do not interact. This is acceptable here because the silica atoms remain frozen during this part of the tutorial.
@@
The number of oxygen atoms from water molecules (i.e. the number of molecules) is calculated by the nO variable. As already discussed in other tutorials, the SHAKE algorithm is used to maintain the shape of the water molecules over time [47, 48].
Note 3.43.
Here, a variable of type ‘atom’ is used. Such variable defines a per-atom property, i.e., it evaluates the specified expression separately for each atom. This is often used to select atoms based on their properties or types.
@@
compute_modify thermo_temp dynamic/dof yes
compute_modify ctH2O dynamic/dof yes
fix mynvt H2O nvt temp 300 300 0.1
fix_modify mynvt temp ctH2O
@@
Here, the fix nvt applies only to the water molecules, so the atoms in the silica remain fixed. The compute_modify command with the dynamic/dof yes option is used for water to account for the fact that the number of molecules is not constant.
@@
The fix gcmc command performs grand canonical Monte Carlo moves to insert, delete, or swap molecules. Here, 100 attempts are made every 100 steps. The mol h2omol keyword specifies the molecule type being inserted/deleted, while shake shak enforces rigid molecular constraints during these moves. With the pressure 100 keyword, a fictitious reservoir with a pressure of 100 atmospheres is used. The tfac_insert option ensures the correct estimate for the temperature of the inserted water molecules by taking into account the internal degrees of freedom.
Note 3.44.
At a pressure of $p=100\,\text{bar}$, the chemical potential of water vapor at $T=300\,\text{K}$ can be calculated using as $\mu=\mu_{0}+RT\ln(\frac{p}{p_{0}}),$ where $\mu_{0}$ is the standard chemical potential (typically taken at a pressure $p_{0}=1\,\text{bar}$), $R=8.314\,\text{J/mol·K}$ is the gas constant, $T=300\,\text{K}$ is the temperature.
@@
The f_ keywords extract the Monte Carlo move statistics which is computed (and can be extracted) by the fix gcmc command.
@@
Note 3.45.
When using the pressure argument, LAMMPS ignores the value of the chemical potential (here $\mu=-0.5\,\text{eV}$, which corresponds roughly to ambient conditions, i.e. to a relative humidity $\text{RH}\approx 50\,\%$ [57].) The large pressure value of 100 bars was chosen to ensure that some successful insertions of molecules would occur during the short duration of this simulation.
Running this simulation using LAMMPS, one can see that after a few GCMC steps, the number of molecules starts increasing. Once the crack is filled with water molecules, the total number of molecules reaches a plateau (Figs. 38-39). The final number of molecules depends on the imposed pressure, temperature, and the interaction between water and silica (i.e. its hydrophilicity). Note that GCMC simulations of such dense phases are usually slow to converge due to the very low probability of successfully inserting a molecule. Here, the short simulation duration was made possible by the use of a high pressure.
@@ 3.7.1 Method 1: Free sampling
The most direct way to estimate a free energy profile is to sample the Boltzmann distribution using a classical (i.e. unbiased) molecular dynamics simulation, and compute relative Gibbs free energies from the relative probabilities of states using
$\Delta G=-RT\ln(\rho/\rho_{0}),$ (3)
where $\Delta G$ is the free energy difference, $R$ is the gas constant, $T$ is the temperature, $\rho$ is the density, and $\rho_{0}$ is a reference density. As an illustration, let us apply this method to a simple configuration that consists of a particles in a box in the presence of a position-dependent repulsive force that makes the center of the box a less favorable area to explore.
@@
To begin this tutorial, select «Start Tutorial 7» from the «Tutorials» menu of LAMMPS–GUI and follow the instructions. The editor should display the following content corresponding to free-sampling.lmp:
@@
Here, we begin by defining variables for the Lennard-Jones interaction $\sigma$ and $\epsilon$ and for the repulsive potential $U$, which are $U_{0}$, $\delta$, and $x_{0}$ [see Eqs. (4-5) below]. The cut-off value of $2^{1/6}\sigma=3.822$ was chosen to create a Weeks-Chandler-Andersen (WCA) potential, which is a truncated and purely repulsive LJ potential [64]. The potential is also shifted to be equal to 0 at the cut-off using the pair_modify command. The unit system is real, in which energy is in kcal/mol, distance in Ångströms, or time in femtosecond, has been chosen for practical reasons: the WHAM algorithm used in the second part of the tutorial automatically assumes the energy to be in kcal/mol.
Note 3.46.
The syntax \$(...), where a dollar sign is followed by parentheses, allows you to evaluate a numeric formula immediately, without having to assign it to a named variable first.
@@
Let us define the simulation box and randomly add atoms by addying the following lines to free-sampling.lmp:
@@
Note 3.47.
In the pair_coeff command, the first two asterisks * * indicate that the parameters apply to all atom types in the simulation.
@@
Fig. 41 shows the potential $U$ and force $F$ along the $x$-axis. With $U_{0}=1.5\epsilon=0.36\,\text{kcal/mol}$, $U_{0}$ is of the same order of magnitude as the thermal energy $k_{\text{B}}T=0.24\,\text{kcal/mol}$, where $k_{\text{B}}=0.002\,\text{kcal/mol/K}$ is the Boltzmann constant and $T=119.8\,\text{K}$ is the temperature used in this simulation. Under these conditions, particles are expected to frequently overcome the energy barrier due to thermal agitation.
Figure 41: Potential $U$ given in Eq. (4) (a) and force $F$ given in Eq. (5) (b) as functions of the coordinate $x$. Here, $U_{0}=0.36\leavevmode\nobreak\ \text{kcal/mol}$, $\delta=1.0\leavevmode\nobreak\ \text{\AA {}}$, and $x_{0}=10\leavevmode\nobreak\ \text{\AA {}}$.
We impose the force $F(x)$ to the atoms in the simulation using the fix addforce command. Add the following lines to free-sampling.lmp:
@@
Next, we use the Newtonian equations of motion with a Langevin thermostat by combining the fix nve with a fix langevin command:
@@
Note 3.48.
LAMMPS documentation suggests using damping constants for thermostats that are approximately 100 times the timestep value. In this case, a value of 500 is used, resulting in a relatively weak coupling to the thermostat.
Figure 42: Evolution of the number of atoms $n_{\text{center}}$ in the central region mymes as a function of time $t$ during equilibration. The dark line is $n_{\text{center}}=22\exp(-t/160)+5$ and serves as a guide for the eyes. Here, $U_{0}=0.36\leavevmode\nobreak\ \text{kcal/mol}$, $\delta=1.0\leavevmode\nobreak\ \text{\AA {}}$, and $x_{0}=10\leavevmode\nobreak\ \text{\AA {}}$.
To ensure that the equilibration time is sufficient, we will track the evolution of the number of atoms in the central - energetically unfavorable - region, defined below under the name mymes, using the n_center variable:
@@
dump viz all image 5000 myimage-*.ppm type type &
@@
A dump image command was also added for system visualization. The other commands should also be familiar from previous tutorials.
Finally, let us perform an equilibration of 50000 steps, using a timestep of $2\,\text{fs}$, corresponding to a total duration of $100\,\text{ps}$:
@@
Run the simulation with LAMMPS. The number of atoms in the central region, $n_{\mathrm{center}}$, reaches its equilibrium value after approximately $40\,\text{ps}$ (Fig. 42). A snapshot of the equilibrated system is shown in Fig. 43.
@@
Once the system is equilibrated, we will record the density profile of the atoms along the $x$-axis using the ave/chunk command. Add the following line to free-sampling.lmp:
@@
Here, the chunk/atom command discretizes the simulation domain into spatial bins of size 2 Å along the $x$ direction, and the ave/chunk command outputs the number density of atoms within each bin to the file free-sampling.dat. The step count is reset to 0 using reset_timestep to synchronize it with the output times of fix density/number. Run the simulation using LAMMPS.
@@
Figure 43: Snapshot of the system simulated during the free sampling step of Tutorial 7. The atoms density is the lowest in the central part of the box, mymes. Here, $U_{0}=0.36\leavevmode\nobreak\ \text{kcal/mol}$, $\delta=1.0\leavevmode\nobreak\ \text{\AA {}}$, and $x_{0}=10\leavevmode\nobreak\ \text{\AA {}}$.
Once the simulation is complete, the density profile from free-sampling.dat shows that the density in the center of the box is about two orders of magnitude lower than inside the reservoir (Fig. 44 a). Next, we plot $RT\ln(\rho/\rho_{\mathrm{bulk}})$, where $\rho/\rho_{\mathrm{bulk}}$ is the the density ratio, and compare it with the imposed potential $U$ from Eq. (4) (Fig. 44 b). The reference density, $\rho_{\text{bulk}}=0.0009\leavevmode\nobreak\ \text{\AA {}}^{-3}$, was estimated by measuring the density of the reservoir from the density profiles. The agreement between the MD results and the imposed energy profile is excellent, despite some noise in the central part, where fewer data points are available due to the repulsive potential.
@@
Increasing the value of $U_{0}$ reduces the average number of atoms in the central region, making it difficult to achieve a high-resolution free energy profile within reasonable simulation times. For example, running the same simulation with $U_{0}=10\epsilon$, corresponding to $U_{0}\approx 10k_{\text{B}}T$, results in no atoms exploring the central part of the simulation box during the simulation. In such a case, employing an enhanced sampling method is recommended, as done in the next section.
@@ 3.7.2 Method 2: Umbrella sampling
Figure 44: a) Fluid density, $\rho$, along the $x$ direction. b) Potential, $U$, as a function of $x$ measured using free sampling (blue disks) compared to the imposed potential given in Eq. (4) (dark line). Here, $U_{0}=0.36\leavevmode\nobreak\ \text{kcal/mol}$, $\delta=1.0\leavevmode\nobreak\ \text{\AA {}}$, $x_{0}=10\leavevmode\nobreak\ \text{\AA {}}$, and the measured reference density in the reservoir is $\rho_{\text{bulk}}=0.0009\leavevmode\nobreak\ \text{\AA {}}^{-3}$.
@@
dump viz all image 5000 myimage-*.ppm type type &
@@
So far, our code resembles that of Method 1, except for the additional particle of type 2. Particles of types 1 and 2 are identical. However, the particle of type 2 will also be exposed to the biasing potential $V$, which forces it to explore the central part of the box (Fig. 45), thus justifying the definition of two atom types.
Figure 45: Snapshot of the system simulated during the umbrella sampling step of Tutorial 7, showing type-1 atoms in cyan and the type-2 atom in red. Only the type-2 atom explores the central part of the box, mymes, due to the additional biasing potential $V$. Parameters are $U_{0}=2.38\leavevmode\nobreak\ \text{kcal/mol}$, $\delta=1.0\leavevmode\nobreak\ \text{\AA {}}$, and $x_{0}=10\leavevmode\nobreak\ \text{\AA {}}$.
@@
The definition of a variable of loop style serves the same purpose as in (Tutorial 5), and we highlight here the particular utility of using its value to distinguish the files written by the fix ave_time command for the different bias potentials. The spring command imposes the additional harmonic potential $V$ with the previously defined spring constant $k$ to the atoms in the group topull. The center of the harmonic potential, $x_{\text{des}}$, successively takes values from $28\,\text{\AA {}}$ to $28\,\text{\AA {}}$. For each value of $x_{\text{des}}$, an equilibration step of 40 ps is performed, followed by a step of 400 ps during which the position of the particle of type 2 along the $x$-axis, $x_{\text{ave}}$, is saved in data files named umbrella-sampling.i.dat, where $i$ ranges from 1 to 15. Run the umbrella-sampling.lmp file using LAMMPS.
Note 3.49.
@@
Figure 46: The potential, $U$, as a function of $x$, measured using umbrella sampling during Tutorial 7 (blue disks), is compared to the imposed potential given in Eq. (4) (dark line). Parameters are $U_{0}=2.38\leavevmode\nobreak\ \text{kcal/mol}$, $\delta=1.0\leavevmode\nobreak\ \text{\AA {}}$, and $x_{0}=10\leavevmode\nobreak\ \text{\AA {}}$.